OpenStudy (anonymous):

please help find an equation for the circle center (-3,-7) radius 1

OpenStudy (akashdeepdeb):



OpenStudy (anonymous):

thanks guys! @AkashdeepDeb so it would just basically be (x-3)^2+(y-7)^2=1

OpenStudy (anonymous):

yes

OpenStudy (gorv):

@lolasally plzz click on dest response if u satisfied by @akashdeepdeb 's explanation

OpenStudy (gorv):

by click best reponse to his ans u will rewars him a medal

OpenStudy (akashdeepdeb):

@lolsally No. The point you have is (-3,-7) and the equation is: \((x - x_1)^2 + (y-y_!)^2 = r^2\) Do you get your error? Be careful with integers! ;-)